### v7

# The defect of weak approximation for a reductive group over a global field

Abstract.
We compute the defect of weak approximation for a reductive group $G$ over a global field $K$ in terms of the algebraic fundamental group of $G$.
Key words and phrases:
Weak approximation, Galois cohomology, reductive groups, global fields
1991 Mathematics Subject Classification:
11E72, 20G10, 20G30
The author was partially supported by the Israel Science Foundation (grant 1030/22).
1. Introduction
$1.$
Let $K$ be a global field (a number field or a global function field), and let $K^{s}$ be a fixed separable closure of $K$. Let $G$ be a reductive group over $K$ (we follow the convention of SGA3, where reductive groups are assumed to be connected). Let $\mathcal{V}_{K}$ denote the set of places of $K$, and let $S\subset\mathcal{V}_{K}$ be a finite set of places. Consider the group
$G(K_{S})\coloneqq\prod_{v\in S}G(K_{v})$
where $K_{v}$ denotes the completion of $K$ at $v$. The group $G(K)$ embeds diagonally into $G(K_{S})$. One says that $G$ has the weak approximation property in $S$ (for short (WAS)) if $G(K)$ is dense in $G(K_{S})$. One says that $G$ has the weak approximation property if it has the weak approximation property in $S$ for any finite subset $S\subset\mathcal{V}_{K}$.
Without assuming that $G$ has the weak approximation property in $S$, let $\overline{G(K)}_{S}$ denote the closure of $G(K)$ in $G(K_{S})$. Sansuc [17, §3] showed (in the number field case) that the subgroup $\overline{G(K)}_{S}$ is normal in $G(K_{S})$ and that the quotient group
$A_{S}(G)\coloneqq G(K_{S})/\overline{G(K)}_{S}$
is a finite abelian group. Sansuc’s argument extends to the function field case. We say that $A_{S}(G)$ is the defect of weak approximation for $G$ in $S$.
1.2.
In the number field case, Sansuc computed $A_{S}(G)$ when $G$ is semisimple, or, more generally, when $G$ admits a special covering, that is, $G$ fits into a short exact sequence of special kind
$1\to B\to G^{\prime}\to G\to 1.$ (1)
Here $G^{\prime}$ is the product of a simply connected semisimple $K$-group and a quasi-trivial $K$-torus, and $B$ is a finite abelian $K$-group. Namely, Sansuc constructed an isomorphism
$A_{S}(G){\ \xrightarrow{\makebox[11.99998pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }\!\!\!\textcyr{Ch}^{1}_{S}(B)\coloneqq\operatorname{coker}\Big{[}H^{1}(K,B)\to\prod_{v\in S}H^{1}(K_{v},B)\Big{]}.$ (2)
Note that there exist reductive $K$-groups not admitting a special covering.
1.3.
[…]
1.4.
In this paper, we compute the defect of weak approximation $A_{S}(G)$ for a reductive $K$-group $G$ in terms of the algebraic fundamental group ${\pi_{1}^{\rm alg}}(G)$ introduced in [2, Section 1] (and also by Merkurjev [11] and Colliot-Thélène [6]). Let $G^{\rm ss}=[G,G]$ denote the derived group of $G$, which is semisimple, and let $G^{\rm sc}$ denote the universal cover of $G^{\rm ss}$, which is simply connected; see [1, Proposition (2:24)(ii)] or [7, Corollary A.4.11]. Consider the composite homomorphism
$\rho\colon G^{\rm sc}\twoheadrightarrow G^{\rm ss}\hookrightarrow G,$
which is in general is neither injective nor surjective. For a maximal torus $T\subseteq G$, we denote
$T^{\rm sc}=\rho^{-1}(T)\subseteq G^{\rm sc}.$
Following [2], we consider the algebraic fundamental group of $G$ defined by
${\pi_{1}^{\rm alg}}(G)={{\sf X}}_{*}(T)/\rho_{*}{{\sf X}}_{*}(T^{\rm sc})$
where $\sf X$$_{*}$ denotes the cocharacter group over $K^{s}$. The absolute Galois group $\operatorname{Gal}(K^{s}/K)$ naturally acts on ${\pi_{1}^{\rm alg}}(G)$, and the Galois module ${\pi_{1}^{\rm alg}}(G)$ is well defined (does not depend on the choice of $T$ up to a transitive system of isomorphisms); see [2, Lemma 1.2]. Note that when $G=T$ is a torus, we have ${\pi_{1}^{\rm alg}}(G)={{\sf X}}_{*}(T)$.
Write $M={\pi_{1}^{\rm alg}}(G)$. Let $T\subseteq G$ be a maximal torus. Choose a finite Galois extension $L/K$ in $K^{s}$ splitting $T$. Set $\Gamma=\operatorname{Gal}(L/K)$, the Galois group of $L$ over $K$. Then the finite group $\Gamma$ naturally acts on $M$. Moreover, $\Gamma$ naturally acts on $L$ and on its set of places $\mathcal{V}_{L}$. We compute $A_{S}(G)$ in terms of the $\Gamma$-module $M$ and the $\Gamma$-set $\mathcal{V}_{L}$.
1.5.
Following Tate [19], we consider the group of finite formal linear combinations
$M[\mathcal{V}_{L}]=\Bigg{\{}\sum_{w\in\mathcal{V}_{L}}m_{w}\cdot w\ \big{|}\ m_{w}\in M\Bigg{\}}$
and its subgroup
$M[\mathcal{V}_{L}]_{0}=\Big{\{}\sum m_{w}\cdot w\in M[\mathcal{V}_{L}]\ \big{|}\ \sum m_{w}=0\Big{\}}.$
Then $\Gamma$ naturally acts on the groups $M[\mathcal{V}_{L}]$ and $M[\mathcal{V}_{L}]_{0}$. We prove the following theorem, which is our main result:
Theorem 1.6 (Theorem 4.6).
$A_{S}(G)\cong\operatorname{coker}\Big{[}H_{1}(\Gamma,M[\mathcal{V}_{L}]_{0})\to\bigoplus_{v\in S}H_{1}({\Gamma\kern-0.8pt_{w}},M)\Big{]}$
where $H_{1}$ denotes group homology. Here for $v\in\mathcal{V}_{K}$, we choose a place $w=w(v)\in\mathcal{V}_{L}$ over $v$, and we denote by $\Gamma\kern-0.8pt_{w}$} the corresponding decomposition group (the stabilizer of $w$ in $\Gamma$).
Theorem 1.6 describes $A_{S}(G)$ in terms of the finite groups $H_{1}({\Gamma\kern-0.8pt_{w}},M)$ and the infinite group $H_{1}(\Gamma,M[\mathcal{V}_{L}]_{0}$. In Section 4 we prove Theorem 4.10, a version of Theorem 1.6, which describes $A_{S}(G)$ in terms of finite groups only, which is good for computer computations.
The plan of the rest of the paper is as follows. In Section 2, for a reductive group $G$ over $K$ splitting over a finite Galois extension $L/K$, we introduce the notion of an $L/K$-free resolution of $G$, and we show that $G$ admits an $L/K$-free resolution. In Section 3 we compute $\textcyr{Ch}^{1}_{S}(B)$ for a $K$-torus $B$ in terms of ${\pi_{1}^{\rm alg}}(B)={{\sf X}}_{*}(B)$. In Section 4 we prove Theorems 4.6 and 4.10 which are our main results. In Section 5 we consider an example over a global function field $K$: we compute $A_{S}(G)$ for a certain $K$-torus $G$.
The arXiv version of this paper contains Appendix A with a listing of a Magma program computing $A_{S}(G)$. Our input is: the effective Galois group $\Gamma$ acting on the algebraic fundamental group $M={\pi_{1}^{\rm alg}}(G)$, the non-cyclic decomposition groups $\Gamma\kern-0.8pt_{w}$ for $v\in S$, and the non-cyclic decomposition groups $\Gamma\kern-0.8pt_{w}$ for $v$ in the complement $S^{\complement}$ of $S$. We compute $A_{S}(G)$ using Theorem 4.10.
Acknowledgements:  The author is grateful to Dipendra Prasad for his suggestion to compute the defect of weak approximation for a reductive group. We thank Will Sawin for his answer [18] to a MathOverflow question of the author, and we thank Boris Kunyavskiĭ for helpful discussions and email correspondence. We thank Willem A. de Graaf for writing the Magma function ComplementIntMat, a Magma version of the GAP function with the same name. This paper was partly written during the author’s visit to Max Planck Institute for Mathematics, Bonn, and the author is grateful to the institute for the hospitality, support, and excellent working conditions.
2. Quasi-trivial groups and $L/K$-free groups
2.1.
A torus $T$ over a field $K$ is called quasi-trivial if its cocharacter group ${{\sf X}}_{*}(T)$ admits a $\operatorname{Gal}(K^{s}/K)$-stable basis. Then by Shapiro’s lemma and Hilbert’s Theorem 90 we have $H^{1}(K,T)=1$. If $K$ is a global field, then $T$ has the weak approximation property (because a quasi-trivial torus is a $K$-rational variety).
We need the following two known theorems.
Theorem 2.2.
Let $G$ be a simply connected semisimple group over a global field $K$, and let $\mathcal{V}_{f}(K)$ and $\mathcal{V}_{\infty}(K)$ denote the sets of finite (non-archimedean) and infinite (archimedean) places of $K$, respectively. Then:
  1. (i)
    For all $v\in\mathcal{V}_{f}(K)$ we have $H^{1}(K_{v},G)=1$.
  2. (ii)
    The localization map
    $\eta\colon\,H^{1}(K,G)\to\prod_{v\in\mathcal{V}(K)}H^{1}(K_{v},G)=\!\!\prod_{v\in\mathcal{V}_{\infty}(K)}H^{1}(K_{v},G)$
    is bijective.
Proof.
Assertion (i) is a theorem of Kneser and of Bruhat and Tits; see Platonov and Rapinchuk [15, Theorem 6.4, p. 284] for the number field case, and Bruhat and Tits [4, Theorem 4.7(ii)] for the general case. Assertion (ii) is the celebrated Hasse principle of Kneser, Harder, and Chernousov; see Platonov and Rapinchuk [15, Theorem 6.6, p. 286] for the number field case, and Harder [8] for the function field case. ∎
Theorem 2.3.
Let $G$ be a simply connected semisimple group over a global field $K$. Then $G$ has the weak approximation property, that is, for any finite subset $S\subset\mathcal{V}(K)$, the group $G(K)$ is dense in $G(K_{S})\coloneqq\prod_{v\in S}G(K_{v})$.
Indeed, $G$ has the weak approximation property because it has the strong approximation property. This was proved by Platonov in characteristic 0 (see Platonov and Rapinchuk [15, Theorem 7.12, p. 427]), and by G. Prasad [16] in positive characteristic.
Definition 2.4 (Colliot-Thélène [6, Definition 2.1]).
A reductive group $G$ over a field $K$ is quasi-trivial if its commutator subgroup $G^{\rm ss}$ is simply connected and the quotient torus $G^{\rm tor}\coloneqq G/G^{\rm tor}$ is a quasi-trivial torus. In other words, $G$ is quasi-trivial if it fits into the exact sequence
$1\to G^{\rm sc}\to G\to G^{\rm tor}\to 1,$ (3)
where $G^{\rm sc}$ is a simply connected semisimple group, and $G^{\rm tor}$ is a quasi-trivial torus.
Proposition 2.5 (Colliot-Thélène [6, Proposition 9.2(i,iii)] in the number field case).
Let $G$ be quasi-trivial group over a global field $K$. Then:
  1. (i)
    For any $v\in\mathcal{V}_{f}(K)$ we have $H^{1}(K_{v},G)=1$.
  2. (ii)
    The localization map
    ${\rm loc}_{\infty}\colon\,H^{1}(K,G)\to\!\!\prod_{v\in\mathcal{V}_{\infty}(K)}\!\!\!\!H^{1}(K_{v},G)$
    is bijective.
Proof.
Concerning (i), from (3) we obtain a cohomology exact sequence
$H^{1}(K_{v},G^{\rm sc})\to H^{1}(K_{v},G)\to H^{1}(K_{v},G^{\rm tor})$
where $H^{1}(K_{v},G^{\rm tor})$ is trivial because $G^{\rm tor}$ is a quasi-trivial torus, and $H^{1}(K_{v},G^{\rm sc})$ is trivial by Theorem 2.2(i). Thus $H^{1}(K_{v},G)$ is trivial.
Concerning (ii), for the case of a number field see [6, Proposition 9.2(iii)]. In the case of a function field, from (3) we obtain a cohomology exact sequence
$H^{1}(K,G^{\rm sc})\to H^{1}(K,G)\to H^{1}(K,G^{\rm tor})$
where $H^{1}(K,G^{\rm tor})$ is trivial because $G^{\rm tor}$ is a quasi-trivial torus, and $H^{1}(K,G^{\rm sc})$ is trivial in the function field case by Theorem 2.2(ii). Thus $H^{1}(K,G)$ is trivial, as desired. ∎
Proposition 2.6 (Colliot-Thélène [6, Proposition 9.2(ii)] in the number field case).
Let $G$ be a quasi-trivial group over a global field $K$. Then $G$ has the weak approximation property, that is, for any finite subset $S\subset\mathcal{V}(K)$, the group $G(K)$ is dense in $G(K_{S})$.
Proof.
The proof of [6, Proposition 9.2(ii)] for the group $A(G)$ in the number field case immediately extends to $A_{S}(G)$ and to the function field case. ∎
Definition 2.7.
Let $T$ be a torus over a field $K$, and let $L/K$ be a finite Galois extension. We say that $T$ is $L/K$-free if $T$ splits over $L$ and the ${\mathds{Z}}[\operatorname{Gal}(L/K)]$-module ${{\sf X}}_{*}(T)$ is free. Alternatively, $T$ is $L/K$-free if it is isomorphic to $(R_{L/K}{{\mathds{G}}_{{\rm m},L}})^{n_{T}}$ for some integer $n_{T}\geq 0$, where ${\mathds{G}}_{{\rm m},L}$} is the multiplicative group over $L$ and $R_{L/K}$ denotes the Weil restriction of scalars.
Observe that any $L/K$-free torus is quasi-trivial.
Definition 2.8.
Let $G$ be a reductive $K$-group with derived group $G^{\rm ss}=[G,G]$. Write $G^{\rm tor}=G/G^{\rm ss}$. We say that $G$ is $L/K$-free if $G^{\rm ss}$ is simply connected and the $K$-torus $G^{\rm tor}$ is $L/K$-free.
Observe that any $L/K$-free reductive $K$-group is quasi-trivial.
Definition 2.9.
Let $G$ be a reductive $K$-group. An $L/K$-free resolution of $G$ is a short exact sequence of $K$-groups
$1\to B\to G^{\prime}\to G\to 1$ (4)
where $G^{\prime}$ is an $L/K$-free reductive $K$-group and $B$ is a $K$-torus.
Proposition 2.10.
Any reductive $K$-group $G$ that splits over a finite Galois extension $L/K$, admits an $L/K$-free resolution.
Proof.
We follow the proof of [6, Proposition-Definition 3.1]. Since $G$ splits over $L$, it has a maximal torus $T$ that splits over $L$. Let $Z$ denote the identity component of the center of $G$ (that is, the radical of $G$); this is a $K$-torus splitting over $L$, because it is a subtorus of the torus $T$ splitting over $L$. There exists a surjective homomorphism of $K$-tori $Q\to Z$ with $Q$ being $L/K$-free. Let $T^{\rm sc}\subset G^{\rm sc}$ denote the preimage of $T$ in $G^{\rm sc}$. Then $T^{\rm sc}$ is isogenous to the $K$-torus $T^{\rm ss}\coloneqq T\cap G^{\rm ss}$. We see that both $T^{\rm ss}$ and $T^{\rm sc}$ split over $L$. We have a natural surjective $K$-homomorphism $G^{\rm sc}\times Q\to G$. Let $Z^{\prime}$ denote the kernel of this homomorphism; then $Z^{\prime}$ is the kernel of the surjective homomorphism $T^{\rm sc}\times Q\to T$, and therefore, $Z^{\prime}$ is a group of multiplicative type over $K$ (not necessary smooth). It follows that we have a short exact sequence of character groups
$0\to{{\sf X}}^{*}(T)\to{{\sf X}}^{*}(T^{\rm sc})\oplus{{\sf X}}^{*}(Q)\to{{\sf X}}^{*}(Z^{\prime})\to 0;$
see Milne [12, Theorem 12.9]. We see that the absolute Galois group $\operatorname{Gal}(K^{s}/K)$ acts on the character group ${{\sf X}}^{*}(Z^{\prime})$ via $\operatorname{Gal}(L/K)$.
By Lemma 2.11 below, there exists a short exact sequence
$1\to Z^{\prime}\to B\to F\to 1,$
where $B$ is a $K$-torus that splits over $L$, and $F$ is an $L/K$-free $K$-torus. The diagonal map defines an embedding of $Z^{\prime}$ into the product $(G^{\rm sc}\times Q)\times B$ with central image. Let $G^{\prime}$ be the quotient of the reductive group $(G^{\rm sc}\times Q)\times B$ by $Z^{\prime}$. Then $G^{\prime}$ is a reductive $K$-group, which fits into the following commutative diagram with exact rows and columns (where we replace the given arrow $Z^{\prime}\to B$ by its inverse):
[…]
The quotient of the group $G^{\prime}$ by the normal subgroup $G^{\rm sc}\times 1\subset G^{\rm sc}\times Q\subset G^{\prime}$ is a $K$-group extension of the $K$-torus $F$ by the $K$-torus $Q$. By [6, Section 0.7] such a $K$-group is a $K$-torus. Since any extension of an $L/K$-free torus by an $L/K$-free torus is a split extension (as one can see on the level of cocharacter groups), we conclude that the quotient of $G^{\prime}$ by $G^{\rm sc}$ is an $L/K$-free torus. The derived group $G^{\prime\,{\rm ss}}$ can be identified with $G^{\rm sc}$, and the group $G^{\prime\,{\rm tor}}$ is an $L/K$-free torus. Thus the reductive group $G^{\prime}$ is $L/K$-free, and the middle row of the diagram is a desired $L/K$-free resolution of $G$. ∎
Lemma 2.11.
Let $\Gamma$ be a finite group and $M$ be a finitely generated $\Gamma$-module. Then there exists a resolution
$0\to M^{-1}\to M^{0}\to M\to 0$
where $M^{0}$ is a finitely generated $\mathds{Z}$}-torsion-free $\Gamma$-module and $M^{-1}$ is a free $\Gamma$-module.
Proof.
See [3, Lemma 4.1.1], or Milne and Shih [13, Proof of Lemma 3.2]. ∎
Theorem 2.12.
Let $G$ be a reductive group over a global field $K$. Consider a short exact sequence
$1\to B\to G^{\prime}\to G\to 1$ (5)
where $G^{\prime}$ is a quasi-trivial $K$-group and $B\subset G^{\prime}$ is a smooth central $K$-subgroup. Let $S$ be a finite set of places of $K$. Then the closure $\overline{G(K)}_{S}$ of $G(K)$ in $G(K_{S})$ is a normal subgroup of finite index, and the connecting homomorphism $G(K_{S})\to H^{1}(K_{S},B)$ induces an isomorphism
$A_{S}(G){\ \xrightarrow{\makebox[12.26656pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }\!\!\!\textcyr{Ch}^{1}_{S}(B).$ (6)
Proof.
Sansuc [17, Theorem 3.3] considers a resolution (5) in the case when $K$ is a number field and $B$ is a finite abelian $K$-group. Moreover, he assumes that the resolution (5) is a special covering. In this case, Sansuc shows that (5) induces an isomorphism (6). Moreover, Colliot-Thélène [6, Theorem 9.4(i)] considers a flasque resolution (5) of $G$, when $K$ is a number field. Then $B$ is a flasque torus. Colliot-Thélène shows that a flasque resolution (5) induces an isomorphism (6).
Our Theorem 2.12 generalizes [17, Theorem 3.3] and [6, Theorem 9.4(i)]. The proof of Sansuc [17] (inspired by Kneser [10]) immediately generalizes to our case. ∎
3. Computing $\textcyr{Ch}^{1}_{S}(B)$ for a torus $B$
Let $B$ be a torus over a global field $K$ with cocharacter group $Y={{\sf X}}_{*}(B)$. Let $L/K$ be a finite Galois extension splitting $T$. We write $\Gamma=\operatorname{Gal}(L/K)$. In this section we compute $\textcyr{Ch}^{1}_{S}(B)$ for a finite set $S$ of places of $K$ in terms of the $\Gamma$-module $Y$. We need two lemmas.
Lemma 3.1 (probably known).
Let $\Delta$} be a finite group and $A$ be a $\Delta$}-module. If $A$ is torsion-free, then
$H^{-1}({\Delta},A)=A_{{\Delta},{\rm Tors}}$
where $A_{\Delta}$ denotes the group of coinvariants of $\Delta$} in $A$, and $A_{{\Delta},{\rm Tors}}\coloneqq(A_{\Delta})_{\rm Tors}$, the torsion subgroup of $A_{\Delta}$.
Proof.
By definition,
$H^{-1}({\Delta},A)=\ker\big{[}{\rm Nm}\colon A_{\Delta}\to A^{\Delta}\big{]}$
where $A^{\Delta}$ denote the group of invariants of $\Delta$} in AA; see [5, Section IV.6]. The group $H^{-1}({\Delta},A)$ is killed by $\Delta$} (see [5, Section IV.6, Corollary 1 of Proposition 8]), whence $H^{-1}({\Delta},A)\subseteq A_{{\Delta},{\rm Tors}}$. Since AA is torsion-free, so is $A^{\Delta}$, whence
$H^{-1}({\Delta},A)=\ker{\rm Nm}\supseteq A_{{\Delta},{\rm Tors}}\kern 0.8pt.$
Thus $H^{-1}({\Delta},A)=A_{{\Delta},{\rm Tors}}$. ∎
Lemma 3.2.
With the notation of the beginning of this section, for any place $v$ of $K$ choose a place $w$ of $L$ over $v$, and denote by ${\Gamma\kern-0.8pt_{w}}\subseteq\Gamma$ the corresponding decomposition group (the stabilizer of $w$ in $\Gamma$). Then we have canonical isomorphisms
$H^{1}(K_{v},B){\ \xrightarrow{\makebox[12.26656pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }Y_{\Gamma\kern-0.8pt_{w}\kern 0.8pt,{\rm Tors}}\kern 0.8pt,\quad\ H^{1}(K,B){\ \xrightarrow{\makebox[12.26656pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }\big{(}Y[\mathcal{V}_{L}]_{0}\big{)}_{\Gamma\kern-0.8pt,\kern 0.8pt{\rm Tors}}\kern 0.8pt.$
Proof.
Since the torus $B$ splits over $L_{w}$, we have $H^{1}(K_{v},B)=H^{1}(L_{w}/K_{v},B)$; see Sansuc [17, (1.9.2)]. We have the Tate-Nakayama isomorphism
$H^{-1}({\Gamma\kern-0.8pt_{w}},Y){\ \xrightarrow{\makebox[11.99998pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }H^{1}({\Gamma\kern-0.8pt_{w}},Y\otimes_{\mathds{Z}}L_{w}^{\times})=H^{1}(L_{w}/K_{v},B);$
see Tate [19, Theorem on page 717]. By Lemma 3.1, we have $H^{-1}({\Gamma\kern-0.8pt_{w}},Y)=Y_{\Gamma\kern-0.8pt_{w}\kern 0.8pt,{\rm Tors}}$, whence $H^{1}(L_{w}/K_{v},B)\cong Y_{\Gamma\kern-0.8pt_{w}\kern 0.8pt,{\rm Tors}}$.
Similarly, we have $H^{1}(K,B)=H^{1}(L/K,B)$; see [17, (1.9.2)]. We have the Tate isomorphism
$H^{-1}\big{(}\Gamma,Y[\mathcal{V}_{L}]_{0}\big{)}{\ \xrightarrow{\makebox[11.99998pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }H^{1}(L/K,B);$
see [19, Theorem on page 717]. Since $Y[\mathcal{V}_{L}]_{0}$ is torsion-free, we conclude by Lemma 3.1 that
$H^{1}(L/K,B)\cong H^{-1}\big{(}\Gamma,Y[\mathcal{V}_{L}]_{0}\big{)}=\big{(}Y[\mathcal{V}_{L}]_{0}\big{)}_{\Gamma\kern-0.8pt,\kern 0.8pt{\rm Tors}}\kern 0.8pt,$
as desired. ∎
$3.$
With the notation of Lemma 3.2, consider the groups
${\Upsilon\!}_{S}=\bigoplus_{v\in S}Y_{\Gamma\kern-0.8pt_{w}\kern 0.8pt,{\rm Tors}}\quad\ \text{and}\quad\ {\Upsilon\!}_{S^{\complement}}=\bigoplus_{v\in S^{\complement}}\!Y_{\Gamma\kern-0.8pt_{w}\kern 0.8pt,{\rm Tors}}\kern 0.8pt.$ (7)
By Lemma 3.2 we have canonical isomorphisms
${\Upsilon\!}_{S}=\bigoplus_{v\in S}H^{1}(K_{v},B)\quad\ \text{and}\quad\ {\Upsilon\!}_{S^{\complement}}=\bigoplus_{v\in S^{\complement}}\!H^{1}(K_{v},B).$
For each $v\in\mathcal{V}_{K}$, consider the natural projection homomorphism
$\pi_{v}\colon Y_{\Gamma\kern-0.8pt_{w}\kern 0.8pt,{\rm Tors}}\to Y_{\Gamma\kern-0.8pt,\kern 0.8pt{\rm Tors}}\kern 0.8pt.$
We have a natural homomorphism
$\pi_{S}\colon\,{\Upsilon\!}_{S}\to Y_{\Gamma\kern-0.8pt,\kern 0.8pt{\rm Tors}}\kern 0.8pt,\ \,\big{[}y_{v}\big{]}_{v\in S}\mapsto\Bigg{[}\,\sum_{v\in S}\pi_{v}(y_{v})\,\Bigg{]}.$
Similarly, we have a natural homomorphism
$\pi_{S^{\complement}}\colon\ {\Upsilon\!}_{S^{\complement}}\to Y_{\Gamma\kern-0.8pt,\kern 0.8pt{\rm Tors}}\kern 0.8pt.$
Theorem 3.4.
There are canonical isomorphisms
$\textcyr{Ch}^{1}_{S}(B)\,\cong{\Upsilon\!}_{S}\kern 0.8pt/\kern 0.8pt\pi_{S}^{-1}(\operatorname{im}\pi_{S^{\complement}})\cong\operatorname{im}\pi_{S}/(\operatorname{im}\pi_{S}\cap\operatorname{im}\pi_{S^{\complement}}).$
Proof.
Consider the following commutative diagram:
${H^{1}(K,B)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\sim}$${H^{-1}\big{(}Y[\mathcal{V}_{L}]_{0}\big{)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\sim}$${\big{(}Y[\mathcal{V}_{L}]_{0}\big{)}_{\Gamma\kern-0.8pt,\kern 0.8pt{\rm Tors}}\thinspace\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{1}(K_{S},B)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\sim}$${\bigoplus\limits_{v\in S}H^{-1}(\Gamma\kern-0.8pt_{w},Y)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\sim}$${\bigoplus\limits_{v\in S}Y_{\Gamma\kern-0.8pt_{w}\kern 0.8pt,{\rm Tors}}\thinspace}$ (8)
in which the rectangle at left comes from Tate [19, Theorem on page 717]. From (8) we obtain a canonical isomorphism
$\textcyr{Ch}^{1}_{S}(B)=\operatorname{coker}\big{[}H^{1}(K,B)\to H^{1}(K_{S},B)\big{]}{\ \xrightarrow{\makebox[11.99998pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }\operatorname{coker}\Big{[}\big{(}Y[\mathcal{V}_{L}]_{0}\big{)}_{\Gamma\kern-0.8pt,\kern 0.8pt{\rm Tors}}\kern-0.8pt\to{\Upsilon\!}_{S}\Big{]}$
where ${\Upsilon\!}_{S}$ is as in (7). From the short exact sequence
$0\to Y[\mathcal{V}_{L}]_{0}\to Y[\mathcal{V}_{L}]\to Y\to 0$
we obtain an exact sequence
$H^{-1}\big{(}\Gamma,Y[\mathcal{V}_{L}]_{0}\big{)}\to\bigoplus\limits_{v\in\mathcal{V}_{K}}H^{-1}(\Gamma\kern-0.8pt_{w},Y)\to H^{-1}(\Gamma,Y),$
which by Lemma 3.1 gives an exact sequence
$\big{(}Y[\mathcal{V}_{L}]_{0}\big{)}_{\Gamma\kern-0.8pt,\kern 0.8pt{\rm Tors}}\ \longrightarrow\ {\Upsilon\!}_{S}\oplus{\Upsilon\!}_{S^{\complement}}\ \xrightarrow{\makebox[35.00005pt]{\scriptsize${\pi_{S}+\pi_{S^{\complement}}}$}}Y_{\Gamma\kern-0.8pt,\kern 0.8pt{\rm Tors}}\kern 0.8pt.$
We see that
$\operatorname{im}\Big{[}\big{(}Y[\mathcal{V}_{L}]_{0}\big{)}_{\Gamma\kern-0.8pt,\kern 0.8pt{\rm Tors}}\kern 0.8pt\to\kern 0.8pt{\Upsilon\!}_{S}\Big{]}\,=\,\pi_{S}^{-1}(\operatorname{im}\pi_{S^{\complement}}),$
which gives the first isomorphism of the theorem. The second isomorphism is obvious. ∎
Corollary 3.5.
Let $S^{\prime}\subset S\subset\mathcal{V}_{K}$ be two finite sets of places of $K$ such that for each place $v\in\Sigma\coloneqq S\smallsetminus S^{\prime}$ and $w\in\mathcal{V}_{L}$ over $v$, there exists a place $v^{\complement}\in S^{\complement}$ and $w^{\complement}\in\mathcal{V}_{L}$ over $v^{\complement}$ with $\Gamma\kern-0.8pt_{w^{\complement}}=\Gamma\kern-0.8pt_{w}$. Then the natural epimorphism $\textcyr{Ch}^{1}_{S}(B)\to\!\!\!\textcyr{Ch}^{1}_{S^{\prime}}(B)$ is an isomorphism.
Proof.
By Theorem 3.4, it suffices to show that for any $v\in\Sigma$, the direct summand $Y_{\Gamma\kern-0.8pt_{w}\kern 0.8pt,{\rm Tors}}$ of ${\Upsilon\!}_{S}$ is contained in $\pi_{S}^{-1}(\operatorname{im}\pi_{S^{\complement}})$. This follows from the equality $\Gamma\kern-0.8pt_{w^{\complement}}=\Gamma\kern-0.8pt_{w}$. ∎
Corollary 3.6 (Sansuc [17]).
Let $S_{\rm nc}\subseteq S$ (resp. $S_{\rm c}\subseteq S$) denote the subset of places with non-cyclic (resp., cyclic) decomposition groups in $\Gamma=\operatorname{Gal}(L/K)$. Then the natural epimorphism $\textcyr{Ch}^{1}_{S}(B)\to\!\!\!\textcyr{Ch}^{1}_{S_{\rm nc}}(B)$ is an isomorphism.
Proof.
Let $v\in S_{\rm c}$ and $w$ be a place of $L$ over $v$. Since the decomposition group $\Gamma\kern-0.8pt_{w}\subset\Gamma$ is cyclic and the set $S$ is finite, by the Chebotarev density theorem, see, for instance, Neukirch [14, Chapter VII, Theorem (13.4)], there exist $v^{\complement}\in S^{\complement}$ and $w^{\complement}\in\mathcal{V}_{L}$ over $v^{\complement}$ with decomposition group $\Gamma\kern-0.8pt_{w^{\complement}}=\Gamma\kern-0.8pt_{w}$, and we conclude by Corollary 3.5. ∎
4. $A_{S}(G)$ in terms of ${{\sf X}}_{*}(B)$ and in terms of $\pi_{1}^{\rm alg}(G)$
4.1.
Let $G$ be a reductive group over a global field $K$. Write $M={\pi_{1}^{\rm alg}}(G)$. Let $T\subseteq G$ be a maximal torus, and let $L/K$ be a finite Galois extension in $K^{s}$ splitting $T$; then $\Gamma\coloneqq\operatorname{Gal}(L/K)$ naturally acts on $M$.
Choose an $L/K$-free resolution (4) as in Definition 2.9, where $B$ is a $K$-torus, and consider the sequence of $\operatorname{Gal}(L/K)$-modules
$0\to Y\to M^{\prime}\to M\to 0$ (9)
where
$Y={\pi_{1}^{\rm alg}}(B)={{\sf X}}_{*}(B)\quad\ \text{and}\quad\ M^{\prime}={\pi_{1}^{\rm alg}}(G^{\prime}).$
This sequence is exact; see [6, Proposition 6.8] or [3, Lemma 6.2.4]. Since $\Gamma$ acts on $Y$ and on $L$, it naturally acts on $\mathcal{V}_{L}$, on $Y[\mathcal{V}_{L}]$, and on $Y[\mathcal{V}_{L}]_{0}$.
Theorem 4.2.
For an $L/K$-free resolution (4) and a finite set of places $S\subset\mathcal{V}_{K}$, there is a canonical isomorphism
$A_{S}(G)\,{\ \xrightarrow{\makebox[12.26656pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }\,\operatorname{coker}\Big{[}\big{(}Y[\mathcal{V}_{L}]_{0}\big{)}_{\Gamma\kern-0.8pt,\kern 0.8pt{\rm Tors}}\to\bigoplus_{v\in S}Y_{\Gamma\kern-0.8pt_{w}\kern 0.8pt,{\rm Tors}}\Big{]}.$
Proof.
From diagram (8) we obtain that
$\!\!\!\textcyr{Ch}^{1}_{S}(B)\coloneqq\operatorname{coker}\Big{[}H^{1}(K,B)\to H^{1}(K_{S},B)\Big{]}\cong\operatorname{coker}\Big{[}\big{(}Y[\mathcal{V}_{L}]_{0}\big{)}_{\Gamma\kern-0.8pt,\kern 0.8pt{\rm Tors}}\to\bigoplus\limits_{v\in S}Y_{\Gamma\kern-0.8pt_{w}\kern 0.8pt,{\rm Tors}}\Big{]}$ ,
and the theorem follows from Theorem 2.12 giving an isomorphism $A_{S}(G){\ \xrightarrow{\makebox[11.99998pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }\!\!\!\textcyr{Ch}^{1}_{S}(B)$. ∎
Theorem 4.3.
For an $L/K$-free resolution (4) and a finite set of places $S\subset\mathcal{V}_{K}$, there is a canonical isomorphism
$A_{S}(G)\,{\ \xrightarrow{\makebox[12.26656pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }\,\operatorname{im}\pi_{S_{\rm nc}}\kern 0.8pt/\big{(}\operatorname{im}\pi_{S_{\rm nc}}\cap\operatorname{im}\pi_{S^{\complement}}\kern 0.8pt\big{)}.$
Proof.
By Theorem 2.12 we have a canonical isomorphism $A_{S}(G){\ \xrightarrow{\makebox[11.99998pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }\!\!\!\textcyr{Ch}^{1}_{S}(B)$, and by Corollary 3.6 we have a canonical isomorphism $\textcyr{Ch}^{1}_{S}(B){\ \xrightarrow{\makebox[11.99998pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }\!\!\!\textcyr{Ch}^{1}_{S_{\rm nc}}(B)$. By Theorem 3.4 we have a canonical isomorphism
$\textcyr{Ch}^{1}_{S_{\rm nc}}(B){\ \xrightarrow{\makebox[11.99998pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }\operatorname{im}\pi_{S_{\rm nc}}/\big{(}\operatorname{im}\pi_{S_{\rm nc}}\cap\operatorname{im}\pi_{(S_{\rm nc})^{\complement}}\kern 0.8pt\big{)}.$
By the Chebotarev density theorem, we have
$\operatorname{im}\pi_{(S_{\rm nc})^{\complement}}=\operatorname{im}\pi_{S^{\complement}}.$
Thus we obtain a desired canonical isomorphism of the theorem. ∎
Corollary 4.4 (Sansuc [17]).
Assume that $G$ splits over a finite Galois extension $L/K$, and let $S_{\rm nc}\subseteq S$ denote the subset of places with non-cyclic decomposition group in $\Gamma=\operatorname{Gal}(L/K)$. Then the natural epimorphism $A_{S}(G)\to A_{S_{\rm nc}}(G)$ is an isomorphism.
Proof.
For an $L/K$-free resolution (4) as in Definition 2.9, where $B$ is a $K$-torus, consider the commutative diagram
${A_{S}(G)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\sim}$${\!\!\!\textcyr{Ch}^{1}_{S}(B)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\sim}$${A_{S_{\rm nc}}(G)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\sim}$${\!\!\!\textcyr{Ch}^{1}_{S_{\rm nc}}(B)}$
in which the horizontal arrows are isomorphisms of Theorem 2.12. By Corollary 3.6 the right-hand vertical arrow is an isomorphism. Thus the left-hand vertical arrow is an isomorphism as well, as desired. ∎
Corollary 4.5 (Sansuc [17, Corollary 3.5]).
  1. (i)
    If $G$ splits over a finite Galois extension $L/K$ and all places in $S$ have cyclic decomposition group in $\operatorname{Gal}(L/K)$, then $G$ has the weak approximation property in $S$.
  2. (ii)
    $G$ has the weak approximation property in $\mathcal{V}_{\infty}(K)$.
Theorem 4.6.
Let $G$ be a reductive group over a global field $K$, and let $M$, $L$, and $\Gamma$ be as in 4.1. Then there is a canonical isomorphism
$A_{S}(G){\ \xrightarrow{\makebox[12.26656pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }\operatorname{coker}\Big{[}H_{1}\big{(}\Gamma,M[\mathcal{V}_{L}]_{0}\big{)}\to\bigoplus_{v\in S}H_{1}(\Gamma\kern-0.8pt_{w},M)\Big{]}$
where $H_{1}$ denotes group homology.
We need a lemma and a corollary.
Lemma 4.7.
Let $\Gamma$ be a finite group, and let $N$ be a finitely generated free ${\mathds{Z}}[\Gamma]$-module. Then
  1. (i)
    $N_{\Gamma\kern-0.8pt,\kern 0.8pt{\rm Tors}}=0$, and
  2. (ii)
    $H_{i}(\Gamma,N)=0$ for all $i\geq 1$.
Proof.
By Shapiro’s lemma we can reduce our lemma to the case $\Gamma=1$ and $M={\mathds{Z}}$, in which assertions (i) and (ii) are obvious. ∎
Corollary 4.8.
Let $L/K$ be a finite Galois extension of global fields with Galois group $\Gamma=\operatorname{Gal}(L/K)$, and let $N$ be a finitely generated free ${\mathds{Z}}[\Gamma]$-module. Then
  1. (i)
    $\big{(}N[\mathcal{V}_{L}]_{0}\big{)}_{\Gamma\kern-0.8pt,\kern 0.8pt{\rm Tors}}=0$, and
  2. (ii)
    $H_{i}(\Gamma,N[\mathcal{V}_{L}]_{0})=0$ for all $i\geq 1$.
Proof.
By [3, Theorem A.1.1], the short exact sequence
$0\to N[\mathcal{V}_{L}]_{0}\to N[\mathcal{V}_{L}]\to N\to 0$
gives rise to a homology exact sequence
$\dots\to H_{2}(\Gamma,N[\mathcal{V}_{L}]_{0})\to H_{2}(\Gamma,$ $N[\mathcal{V}_{L}])\to H_{2}(\Gamma,N)$
$\to H_{1}(\Gamma,N[\mathcal{V}_{L}]_{0})\to$ $H_{1}(\Gamma,N[\mathcal{V}_{L}])\to H_{1}(\Gamma,N)$ (10)
$\to\big{(}N[\mathcal{V}_{L}]_{0}$ $\big{)}_{\Gamma\kern-0.8pt,\kern 0.8pt{\rm Tors}}\to\big{(}N[\mathcal{V}_{L}]\big{)}_{\Gamma\kern-0.8pt,\kern 0.8pt{\rm Tors}}\to N_{\Gamma\kern-0.8pt,\kern 0.8pt{\rm Tors}}$
By Lemma 4.7 we have $H_{i}(\Gamma,N)=0$ for all $i\geq 1$. Moreover, we have canonical isomorphisms
$H_{i}(\Gamma,N[\mathcal{V}_{L}]\cong\bigoplus_{v\in\mathcal{V}_{K}}\!H_{i}(\Gamma\kern-0.8pt_{w},N)=0\ \,\ \text{for all $i\geq 1,$}$
$\big{(}N[\mathcal{V}_{L}]\big{)}_{\Gamma\kern-0.8pt,\kern 0.8pt{\rm Tors}}\cong\bigoplus_{v\in\mathcal{V}_{K}}\!N_{\Gamma\kern-0.8pt_{w}\kern 0.8pt,{\rm Tors}}=0\$
by Shapiro’s lemma and Lemma 4.7 (because the ${\mathds{Z}}[\Gamma]$-free module $N$ is ${\mathds{Z}}[\Gamma\kern-0.8pt_{w}]$-free). Now the corollary follows from the exactness of (10). ∎
Proof of Theorem 4.6.
For an $L/K$-free resolution of Definition 2.9, the short exact sequence of $\operatorname{Gal}(L/K)$-modules (9) induces a short exact sequence
$0\to Y[\mathcal{V}_{L}]_{0}\to M^{\prime}[\mathcal{V}_{L}]_{0}\to M[\mathcal{V}_{L}]_{0}\to 0$
and a commutative diagram with exact rows
${0=H_{1}(\Gamma,M^{\prime}[\mathcal{V}_{L}]_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{1}(\Gamma,M[\mathcal{V}_{L}]_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{l_{1}}$${\big{(}Y[V_{L}]_{0}\big{)}_{\Gamma\kern-0.8pt,\kern 0.8pt{\rm Tors}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{l_{0}}$${\big{(}M^{\prime}[V_{L}]_{0}\big{)}_{\Gamma\kern-0.8pt,\kern 0.8pt{\rm Tors}}=0}$${0=\bigoplus\limits_{v\in S}\!H_{1}(\Gamma\kern-0.8pt_{w},M^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bigoplus\limits_{v\in S}\!H_{1}(\Gamma\kern-0.8pt_{w},M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bigoplus\limits_{v\in S}\!Y_{\Gamma\kern-0.8pt_{w}\kern 0.8pt,{\rm Tors}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bigoplus\limits_{v\in S}(M^{\prime})_{\Gamma\kern-0.8pt_{w}\kern 0.8pt,{\rm Tors}}=0}$
In this diagram, the zeros in the bottom row are explained by Lemma 4.7, and the zeros in the top row are explained by Corollary 4.8. The diagram induces an isomorphism $\operatorname{coker}l_{1}{\ \xrightarrow{\makebox[11.99998pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }\operatorname{coker}l_{0}$. By Theorem 4.2 we have an isomorphism $A_{S}(G)\cong\operatorname{coker}l_{0}$, which gives a desired isomorphism $A_{S}(G)\cong\operatorname{coker}l_{1}$. ∎
4.9.
Consider the sets $S$, $S^{\complement}$; then $S\cup S^{\complement}=\mathcal{V}_{K}$. Let $S_{\rm nc}\subseteq S$ and $(S^{\complement})_{\rm nc}\subset S^{\complement}$ be the subsets consisting of all places $v$ with non-cyclic decomposition group $\Gamma\kern-0.8pt_{w}$ where $w$ is a place of $L$ over $v$. Let ${\rm Cyc}(\Gamma)$ denote the set of all cyclic subgroups ${\Delta}\subseteq\Gamma$.
Consider the groups
$\Xi_{S}=\bigoplus_{v\in S}H_{1}(\Gamma\kern-0.8pt_{w},M)\quad\ \text{and}\quad\ \Xi_{S^{\complement}}=\bigoplus_{v\in S^{\complement}}\!H_{1}(\Gamma\kern-0.8pt_{w},M).$
Moreover, consider the groups
$\Xi_{S_{\rm nc}}=\bigoplus_{v\in S_{\rm nc}}H_{1}(\Gamma\kern-0.8pt_{w},M)\quad\ \text{and}\quad\ \Xi_{(S^{\complement})_{\rm nc}}=\bigoplus_{v\in(S^{\complement})_{\rm nc}}\!H_{1}(\Gamma\kern-0.8pt_{w},M).$
Furthermore, consider the group
$\Xi_{\rm Cyc}=\bigoplus_{{\Delta}\in{\rm Cyc}(\Gamma)}H_{1}({\Delta},M).$
For $v\in\mathcal{V}_{K}$ consider the natural corestriction homomorphism
$\tau_{v}\colon H_{1}(\Gamma\kern-0.8pt_{w},M)\to H_{1}(\Gamma,M).$
We have a natural homomorphism
$\tau_{S\phantom{{}^{\complement}}}\colon\,\Xi_{S}\to H_{1}(\Gamma,M),\ \,\big{[}\xi_{v}\big{]}_{v\in S}\mapsto\Bigg{[}\,\sum_{v\in S_{\rm nc}}\tau_{v}(\xi_{v})\,\Bigg{]}.$
Moreover, we have a natural homomorphism
$\tau_{S^{\complement}}\!\colon\,\Xi_{S^{\complement}}\!\to H_{1}(\Gamma,M).$
Furthermore, we have natural homomorphisms
$\tau_{S_{\rm nc}\phantom{{}^{\complement}}\!\!}\colon\,\Xi_{S_{\rm nc}}\!\to H_{1}(\Gamma,M),\ \quad\tau_{(S^{\complement})_{\rm nc}}\colon\,\Xi_{(S^{\complement})_{\rm nc}}\!\to H_{1}(\Gamma,M),\ \quad\tau_{{\rm Cyc}{\phantom{{}^{\complement}}}}\!\!\colon\,\Xi_{\rm Cyc}\to H_{1}(\Gamma,M).$
Theorem 4.10.
There are canonical isomorphisms
$A_{S}(G)\,\cong\,\operatorname{im}\tau_{S_{\rm nc}\phantom{{}^{\complement}}}\!/\big{(}\operatorname{im}\tau_{S_{\rm nc}\phantom{{}^{\complement}}}\!\!\!\cap\operatorname{im}\tau_{S^{\complement}}\kern 0.8pt\big{)}\,\cong\,\operatorname{im}\tau_{S_{\rm nc}\phantom{{}^{\complement}}}\!/\big{(}\operatorname{im}\tau_{S_{\rm nc}\phantom{{}^{\complement}}}\!\!\!\cap(\operatorname{im}\tau_{(S^{\complement})_{\rm nc}}\!+\operatorname{im}\tau_{{\rm Cyc}{\phantom{{}^{\complement}}}}\!\!)\kern 0.8pt\big{)}.$
Proof.
For each $v\in\mathcal{V}_{K}$, the short exact sequence (9) gives rise to a commutative diagram with exact rows
${0=H_{1}(\Gamma\kern-0.8pt_{w},M^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{1}(\Gamma\kern-0.8pt_{w},M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tau_{v}}$${Y_{\Gamma\kern-0.8pt_{w}\kern 0.8pt,{\rm Tors}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{v}}$${(M^{\prime})_{\Gamma\kern-0.8pt_{w}\kern 0.8pt,{\rm Tors}}=0}$${0=H_{1}(\Gamma,M^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{1}(\Gamma,M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Y_{\Gamma\kern-0.8pt,\kern 0.8pt{\rm Tors}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(M^{\prime})_{\Gamma\kern-0.8pt,\kern 0.8pt{\rm Tors}}=0}$
In this diagram, the zeros are explained by Lemma 4.7. Now we obtain the first isomorphism of the theorem from Theorem 4.3 giving isomorphisms
$A_{S}(G)\,{\ \xrightarrow{\makebox[11.99998pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }\,{\Upsilon\!}_{S_{\rm nc}}\kern 0.8pt/\kern 0.8pt\pi_{S_{\rm nc}}^{-1}\big{(}\operatorname{im}\pi_{S^{\complement}}\big{)}\,\cong\,\operatorname{im}\pi_{S_{\rm nc}}\kern 0.8pt/\big{(}\operatorname{im}\pi_{S_{\rm nc}}\cap\operatorname{im}\pi_{S^{\complement}}\kern 0.8pt\big{)},$
and from the isomorphisms
$H_{1}(\Gamma\kern-0.8pt_{w},M){\ \xrightarrow{\makebox[11.99998pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }Y_{\Gamma\kern-0.8pt_{w}\kern 0.8pt,{\rm Tors}}\kern 0.8pt,\quad\ H_{1}(\Gamma,M){\ \xrightarrow{\makebox[11.99998pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }Y_{\Gamma\kern-0.8pt,\kern 0.8pt{\rm Tors}}$
coming from the above diagram. Further, by the Chebotarev density theorem we have
$\operatorname{im}\tau_{S^{\complement}}=\operatorname{im}\tau_{(S^{\complement})_{\rm nc}}\!+\operatorname{im}\tau_{{\rm Cyc}{\phantom{{}^{\complement}}}}\!,$
which gives the second isomorphism of the theorem. ∎
Corollary 4.11.
Assume that there exist a place $v\in S_{\rm nc}$ of $K$ and a place $w$ of $L$ over $v$ such that $\Gamma\kern-0.8pt_{w}=\Gamma$. Then
$A_{S}(G)\cong\operatorname{coker}\tau_{S^{\complement}}.$
Proof.
Indeed, then the homomorphism $\tau_{S_{\rm nc}\phantom{{}^{\complement}}\!\!}$ is surjective. ∎
Corollary 4.12.
Assume that there exist a place $v\in S^{\complement}$ of $K$ and a place $w$ of $L$ over $v$ such that $\Gamma\kern-0.8pt_{w}=\Gamma$. Then
$A_{S}(G)=0.$
Proof.
Indeed, then the homomorphism $\tau_{S^{\complement}}$ is surjective. ∎
5. An example in positive characteristic
We take $K={\mathds{F}}_{p}(t)$, the field of rational functions in one variable $t$ over a finite field ${\mathds{F}}_{p}$ where $p$ is a prime. We assume that $p\equiv-1\pmod{4}$, for instance, $p=3$ or $p=7$, and we take $L={\mathds{F}}_{p}\big{(}\sqrt{t},\sqrt{t^{2}-1}\,\big{)}$. Then $L/K$ is a Galois extension with Galois group $\Gamma=\operatorname{Gal}(L/K)\simeq{\mathds{Z}}/2{\mathds{Z}}\times{\mathds{Z}}/2{\mathds{Z}}$.
Lemma 5.1.
Let $K={\mathds{F}}_{p}(t)$, $L={\mathds{F}}_{p}(\sqrt{t},\sqrt{t^{2}-1})$ where $p\equiv-1\bmod 4$. Then the decomposition groups for $L/K$ at the places corresponding to the irreducible polynomials $t$ and $t+1$ coincide with $\Gamma$, and all the other decomposition groups are cyclic.
Proof.
[…]
The first extension ramifies at $0,\infty$. At $t=0$, the second extension $y^{2}=t^{2}-1$ locally looks like $y^{2}=0-1=-1$, which is a field since $q\equiv-1\bmod 4$. Thus $L_{w}/K_{v}$ is a composite of a ramified quadratic extension and an unramified quadratic extension of non-archimedean local fields, and hence the Galois group $\Gamma\kern-0.8pt_{w}=\operatorname{Gal}(L_{w}/K_{v})$ is isomorphic to the Klein four-group $\Gamma$. Thus $\Gamma\kern-0.8pt_{w}=\Gamma$. We can write the second extension as $\left(\frac{y}{t}\right)^{2}=1-\frac{1}{t^{2}}$, which at $t=\infty$ locally looks like $\left(\frac{y}{t}\right)^{2}=1-\frac{1}{\infty^{2}}=1$, which is split. Thus $L_{w}/K_{v}$ is a quadratic extension, and therefore $\Gamma\kern-0.8pt_{w}=\operatorname{Gal}(L_{w}/K_{v})$ is a cyclic group of order 2.
The second extension ramifies at $t=1$, where the first extension locally looks like $y^{2}=1$, hence splits, and therefore $\Gamma\kern-0.8pt_{w}$ is a cyclic group of order 2. Moreover, it ramifies at t=1t=-1, where the first extension looks like $y^{2}=-1$, hence is a field, because $p\equiv-1\bmod 4$. As above, we see that $\Gamma\kern-0.8pt_{w}=\Gamma$ in this case.
We see that $\Gamma\kern-0.8pt_{w}=\Gamma$ at the places corresponding to the irreducible polynomials $t$ and $t+1$, and $G_{w}$ is cyclic for all other places $v$ of $K$, as desired. ∎
Proposition 5.2.
Let $T=R^{1}_{L/K}{{\mathds{G}}_{\rm m}}\coloneqq\ker\big{[}R_{L/K}{{\mathds{G}}_{\rm m}}\to{{\mathds{G}}_{\rm m}}\big{]}$ be the norm 1 torus, where $L/K={\mathds{F}}_{p}\big{(}\sqrt{t},\sqrt{t^{2}-1}\,\big{)}/{\mathds{F}}_{p}(t)$ with $p\equiv-1\pmod{4}$. Let $S$ be a finite set of places of $K$. If $S$ contains both places of $K={\mathds{F}}_{p}(t)$ corresponding to the polynomials $t$ and $t+1$, then $A_{S}(T)\cong{\mathds{Z}}/2{\mathds{Z}}$. Otherwise we have $A_{S}(T)=0$.
Proof.
This is similar to [17, Example 5.6] in the number field case. Consider the algebraic fundamental group
$M=\pi_{1}^{\rm alg}(T)={{\sf X}}_{*}(T)=\ker\big{[}\kern 0.8pt{\mathds{Z}}[\Gamma]\to{\mathds{Z}}\kern 0.8pt\big{]}.$
We wish to compute $H^{-2}(\Gamma,M)$ and $H^{-2}(\Gamma\kern-0.8pt_{w},M)$. From the short exact sequence
$0\to M\to{\mathds{Z}}[\Gamma]\to{\mathds{Z}}\to 0$
we obtain an exact sequence of Tate cohomology groups
$0=H^{-3}(\Gamma,{\mathds{Z}}[G])\to H^{-3}(\Gamma,{\mathds{Z}})\to H^{-2}(\Gamma,M)\to H^{-2}(\Gamma,{\mathds{Z}}[\Gamma])=0.$
Thus
$H^{-2}(\Gamma,M)\cong H^{-3}(\Gamma,{\mathds{Z}})=H_{2}(\Gamma,{\mathds{Z}}).$
Similarly we obtain that
$H^{-2}(\Gamma\kern-0.8pt_{w},M)\cong H^{-3}(\Gamma\kern-0.8pt_{w},{\mathds{Z}})=H_{2}(\Gamma\kern-0.8pt_{w},{\mathds{Z}}).$
Now $H_{2}(\Gamma,{\mathds{Z}})$ is the Schur multiplier of $\Gamma$, and by a theorem of Schur we have
$H_{2}(\Gamma,{\mathds{Z}})\cong{\mathds{Z}}/2{\mathds{Z}};$
see [9, Corollary 2.2.12]. Thus $H^{-2}(\Gamma,M)\cong{\mathds{Z}}/2{\mathds{Z}}$.
If $\Gamma\kern-0.8pt_{w}=\Gamma$, then, of course,
$H^{-2}(\Gamma\kern-0.8pt_{w},M)=H^{-2}(\Gamma,M)\cong{\mathds{Z}}/2{\mathds{Z}}.$
If $\Gamma\kern-0.8pt_{w}$ is cyclic, then we have
$H^{-2}(\Gamma\kern-0.8pt_{w},M)\cong H^{-3}(\Gamma\kern-0.8pt_{w},{\mathds{Z}})\cong H^{1}(\Gamma\kern-0.8pt_{w},{\mathds{Z}})=\operatorname{Hom}(\Gamma\kern-0.8pt_{w},{\mathds{Z}})=0,$
where $H^{-3}(\Gamma\kern-0.8pt_{w},{\mathds{Z}})\cong H^{1}(\Gamma\kern-0.8pt_{w},{\mathds{Z}})$ by periodicity; see [5, Section IV.8, Theorem 5].
Now the proposition follows from Corollaries 4.11 and 4.12. ∎
Appendix A Computer program
Here is the computer program computing $A_{S}(G)$ mentioned in the Introduction. We assume that we know the effective Galois group $\Gamma$ acting on $M={\pi_{1}^{\rm alg}}(G)$, all non-cyclic decomposition groups (up to conjugacy) $\Gamma\kern-0.8pt_{w}\subset\Gamma$ for $v\in S$ and $w\in\mathcal{V}_{L}$ over $v$, and all non-cyclic decomposition groups $\Gamma\kern-0.8pt_{w}$ for $v\in S^{\complement}$. We use Theorem 4.10.
/* This program computes the defect
[…]
over a global field K.
We assume that we know:
[…]
of order $g$ with multiplication law given
by a martrix mu of dimensions $g$ x $g$ ;
[…]
given by $g$ matrices $m$ x $m$ ;
* the set of all non-cyclic (and maybe some cyclic)
decomposition subgroups G_v of G for v in S ;
* the set of all non-cyclic (and maybe some cyclic)
decomposition subgroups G_v of G
for v in the commplement S^C of C.
We useTheorem 4.10.
ComplementIntMat:= function( full, sub )
/* This is a Magma version of the GAP function
with the same name.
This function was written by Willem A. de Graaf.*/
full:= Matrix( Integers(), full );
H:= HermiteForm( full );
full:= [ ];
for i in [1..NumberOfRows(H)] do
if not IsZero( H[i] )
then Append( ~full, Eltseq( H[i] ) );
end if;
end for;
full:= Matrix( Integers(), full );
n:= NumberOfColumns(full);
m:= NumberOfRows( full );
Vn:=KModule( Rationals(), n );
W:= KModuleWithBasis( [ Vn!Eltseq(full[i]) : i in [1..m] ] );
cfs:= [ Coordinates( W, Vn!Eltseq(sub[i]) ) :
i in [1..NumberOfRows(sub)] ];
A:= Matrix( Integers(), cfs );
S,P,Q:= SmithForm( A );
Qi:= Q^-1;
cfs:= [ Eltseq(Qi[j]) : j in [1..m] ];
vecs:= [ &+[ c[i]*(Vn!Eltseq(full[i])) :
i in [1..m] ] : c in cfs ];
x:= [ ];
for i in [1..NumberOfColumns(S)] do
if S[i][i] ne 0 then
Append( ~x, S[i][i] );
end if;
end for;
comp:= [ ];
for i in [#x+1..#vecs] do Append( ~comp, vecs[i] ); end for;
return Matrix( Integers(), comp), Matrix( Integers(),
[ Eltseq(vecs[i]) : i in [1..#x] ] ),
Vector( Integers(), x );
end function;
//CHECK FOR mu
IsGroup := function($g$, mu)
/* This fuction checks whether the matrix mu
defines a group structure on the set {1,2,…,gg},
and if yes, it computed the vector iota
describing the inverse map.*/
boo:=true;
for g:=1 to $g$ do
boo := boo and (mu[g,1] eq g) and (mu[1,g] eq g);
// Checking the unit condition
end for;
for g1:=1 to $g$ do
for g2:=1 to $g$ do
for g3:=1 to $g$ do
p12:=mu[g1,g2]; p23:=mu[g2,g3];
boo:=boo and (mu[p12,g3] eq mu[g1,p23]);
// Checking the associativity condition
end for;
end for;
end for;
iota:=[];
for g:=1 to $g$ do
boog:=false;
for h:=1 to $g$ do
boo1:=(mu[g][h] eq 1);
if boo1 then
iota[g]:=h;
boog:=true;
break;
end if;
end for;
boo:=boo and boog;
// Checking the existence of an inverse element
end for;
return boo,iota;
end function;
// CHECK FOR N
IsAction:= function($g$,mu,$m$,rho,$n$,N)
boo:=true;
Zv:=RModule(Integers(), $m$);
v0:= Zero(Zv);
vv:= Zero(Zv);
for g:=1 to $g$ do
for i:=1 to $n$ do
for j:=1 to $m$ do
vv[j]:=N[i][j];
end for;
vv:=$v$ * rho[g];
ic:=IsConsistent(N, $v$);
// Checking that rho preserves <N>,
// the group generated by the rows of the matric N
boo:=boo and ic;
end for;
end for;
for g:=1 to $g$ do
for h:= 1 to $g$ do
Dif:=rho[mu[g][h]] - rho[g]*rho[h];
for i:=1 to $m$ do
for j:=1 to $m$ do
vv[j]:=Dif[i][j];
end for;
ic:=IsConsistent(N, $v$); //print ic;
// Checking that rho is a well-defined
// action modulo <N>
boo:=boo and ic;
end for;
end for;
end for;
return boo;
end function;
NullModN:= function(A,N)
AN:=VerticalJoin(A,N);
K:=KernelMatrix(AN);
rA:=NumberOfRows(A);
rK:=NumberOfRows(K);
B:=ExtractBlockRange(K, 1, 1, rK,rA);
Meet:=B*A;
return B, Meet;
// <B> is the preimage of <N> under A
// <Meet> is the intersection of <A> and <N>.
end function;
Minus2CoCycles:= function($g$, mu, iota, $m$, rho, $n$, N)
[…]
matz:=ZeroMatrix(Integers(), $m$, $m$ );
Idm:=ScalarMatrix(Integers(), $m$, 1);
for h:=1 to $g$ do
matz:=rho[h]-Idm;
if h eq 1 then
Dm1:=matz;
else
Dm1:=VerticalJoin(Dm1,matz);
end if;
end for;
Zm2:=NullModN(Dm1,N);
return Zm2;
// <Zm2> is our group of -2 cocycles
end function;
Minus2CoBoundaries:= function($g$, mu, iota, $m$, rho, $n$, N)
[…]
Zv:=RModule(Integers(), $m$);
v0:= Zero(Zv);
Idm:=ScalarMatrix(mm,1);
Bm2:=ZeroMatrix(Integers(), gg*gg*$m$, gg*$m$ );
c0:=[ []:i in [1..gg]];
for g in [1..gg] do
for h in [1..gg] do
c0[g,h]:=v0;
end for;
end for;
matz:= [Matrix(Integers(), Idm) : u in [1..mm] ];
// We have a triple cycle
for g0:=1 to $g$ do
for h0:=1 to $g$ do
for m0:=1 to $m$ do
c:=c0;
c[g0][h0][m0]:=1;
// We consider a 3-dimensional zero array c
// of dimensions $g$ x $g$ x $m$.
// All values c[g,h,m] are zero except for
// c[g0,h0,m0] that equals 1.
// We obtain $g$ * $g$ * $m$ vectors c[g,h]
// in our module M of rank $m$.
// Then we compute the boundary dc
// by the following double loop:
for g:=1 to $g$ do
sum:=v0;
for h:=1 to $g$ do
sum:=sum + c[h,g]*rho[h]
- c[h, mu[iota[h],g]] + c[g,h];
end for;
i1:=m0+mm*(h0-1)+gg*mm*(g0-1);
for m:=1 to $m$ do
j1:=mm*(g-1)+m;
Bm2[i1][j1]:=sum[m];
end for;
end for;
end for;
end for;
end for;
for g:= 1 to $g$ do
if g eq 1 then
ggN:=N;
else
ggN:=DiagonalJoin(ggN,N);
end if;
end for;
Bm2:=VerticalJoin(Bm2,ggN);
// <Bm2> is our group of -2 coboundaries modulo <N>
return ggN, Bm2;
end function;
Zm2FromCyclic:= function($g$, mu, iota, $m$, rho, $n$, N)
Idm:=ScalarMatrix(Integers(),mm,1);
for g:= 1 to $g$ do
Fixg,Meet:=NullModN(Idm-rho[g],N);
// Fixg are the elements of Z^$m$ that are fixed
// by iota[g] modulo <N>
if g eq 1 then
ALDiag:=Fixg;
else
ALDiag:=DiagonalJoin(ALDiag,Fixg);
end if;
end for;
return ALDiag;
// <ALDiag> are the images of (-2)-cocycles for all g.
end function;
AssignmentsH:= function($g$,mu,iota,$m$,rho, $h$,phiH)
// phiH is the embedding H={1..hh} into G={1..gg}
// we compute the restrictions to H of mu, iota, and rho.
muH:=ZeroMatrix(Integers(),$h$,$h$);
iotaH:=[];
psiH:=[];
for i:=1 to $g$ do
psiH[i]:=0;
for j:=1 to $h$ do
if i eq phiH[j] then
psiH[i] := j;
break;
end if;
end for;
end for;
// psiH is the inverse function for phiH
for g:=1 to $h$ do
for h:=1 to $h$ do
i:=mu[phiH[g],phiH[h]];
muH[g][h]:=psiH[i];
if psiH[i] eq 0 then
[…]
end if;
end for;
end for;
// muH is the restriction to H
// of the multiplication law mu in G
for h:=1 to $h$ do
i:=iota[phiH[h]];
iotaH[h]:=psiH[i];
if psiH[i] eq 0 then
[…]
end if;
end for;
// iotaH is the restriction to H
// of the inversion map iota for G
[…]
rhoH:=[ Matrix(Integers(),rho[phiH[u]]) : u in [1..hh] ];
[…]
//rhoH is the restriction to H of the action rho of G
return psiH, muH, iotaH, rhoH;
end function;
Xi:= function($g$,mu,iota,$m$,rho,$n$,N,phi)
[…]
// for S given by phi
// Here phi is a list of emebeddings phi[l]
// of subgroups H[l] into G
nD:=#(phi);
AL:=ZeroMatrix(Integers(), 1,gg*$m$);
if nD eq 0 then
return AL;
else
for l:=1 to nD do
hh:=#(phi[l]);
phiH:=phi[l];
psiH, muH, iotaH, rhoH
:= AssignmentsH($g$,mu,iota,$m$,rho, $h$,phiH);
Z2H:=Minus2CoCycles($h$,muH,iotaH,$m$,rhoH,$n$,N);
rZ2H:=NumberOfRows(Z2H);
A:=ZeroMatrix(Integers(), rZ2H, gg*$m$);
for i:=1 to rZ2H do
for h:=1 to $h$ do
for j:=1 to $m$ do
A[i][(phiH[h]-1)*mm+j]
:= Z2H[i][(h-1)*mm+j];
// We embed each 1-cycle of H into G
end for;
end for;
end for;
AL:=VerticalJoin(AL,A);
end for;
end if;
return AL;
// AL is the list of all 1-cycles of of G
// coming from 1-cycles of H
// for all H in S, where S and the subgroups H in S
// (subgroups of G) are given by phi
end function;
[…]
gg:=4;
mu:=[[1,2,3,4],[2,1,4,3],[3,4,1,2],[4,3,2,1]];
// Kleins four-group
mm:=3;
Idm:=ScalarMatrix(Integers(), mm,1);
rho:= [ Matrix(Integers(),Idm) : u in [1..gg] ];
if $m$ eq 1 then
rho[2]:=[[3]];
rho[3]:=[[5]];
rho[4]:=[[7]];
end if;
if $m$ eq 4 then
rho:= [ Matrix(Integers(),Idm) :
u in [1..gg] ];
for g:=1 to $g$ do
rhr:=ZeroMatrix(Integers(), $g$, $g$);
for h:=1 to $g$ do
rhr[h,mu[g,h]]:=1;
end for;
rho[g]:=rhr;
end for;
end if;
if $m$ eq 3 then
for g:=1 to $g$ do
rhr:=ZeroMatrix(Integers(), $g$, $g$);
for h:=1 to $g$ do
rhr[h,mu[g,h]]:=1;
end for;
U:=Matrix(RationalField(), $4,$
[1,0,0,-1, 0,1,0,-1, 0,0,1,-1, 1,1,1,1]);
rhr:=U*rhr*U^(-1);
rho[g]:=ExtractBlock(rhr,1,1,$m$,$m$);
end for;
end if;
nn:=0;
if $n$ eq 0 then
N:=ZeroMatrix(Integers(), 1, $m$);
end if;
if $n$ eq 1 then
N:=ZeroMatrix(Integers(), $n$,$m$);
for m:=1 to $m$ do
N[1,m]:=1;
end for;
end if;
if $n$ eq 3 then
N:=4*ScalarMatrix(Integers(),$m$, 1);
end if;
if $n$ eq 5 then
N:= ZeroMatrix(Integers(), mm+1,$m$);
for m:=1 to $m$ do
N[m,m]:=1;
N[mm+1,m]:=1;
end for;
end if;
phiNCyc:=[[1,2,3,4]];
phiC:=[[1,2]];
phiS:=[[1,2,3,4],[1,3]];
[…]
boo, iota := IsGroup($g$,mu);
[…]
IsAction($g$,mu,$m$,rho,$n$, N);
[…]
Zm2:=Minus2CoCycles($g$, mu, iota, $m$, rho, $n$, N);
ggN,Bm2:=Minus2CoBoundaries($g$, mu, iota, $m$, rho, $n$, N);
// ggN is DiagonalJoin N $g$ times.
comp, bas,x:=ComplementIntMat(Zm2,Bm2);
[…]
print x = ”, x;
[…]
corresponding to all cyclic decomposition subgroups,
a list of all non-cyclic (and maybe some cyclic)
decomposition subgroups in the complement S^C of S,
and a list of non-cyclic (and maybe some cyclic)
decomposition subgroups in S”;
[…]
print Below phiH is the emgedding
of H=[1..hh] into Gamma=[1..gg].”;
print We compute the restrictions to H of mu, iota, and rho.”;
ALCyclic:=Zm2FromCyclic($g$, mu, iota, $m$, rho, $n$, N);
AL:=VerticalJoin(Bm2,ALCyclic);
ALC:= Xi($g$,mu,iota,$m$,rho,$n$,N,phiC);
ALC:=VerticalJoin(AL,ALC);
// ALC is the vertical Join of Zm2 from all cyclic subgroups,
// Bm2 (all coboundaries and ggN),
// and (-2)-Cocycles for *non-cyclic* decomposition subgroups
// in the *complement* S^C of the set S.
// Here ggN is the Diagonal Join N $g$ times.
// ALC corresponds to
[…]
// in Theorem 4.10.
ALS:= Xi($g$,mu,iota,$m$,rho,$n$,N,phiS);
ALS:=VerticalJoin(ALS,Bm2);
// ALS is the vertical Join of Zm2 from all
// *non-cyclic* decomposition subgroups in the set S
// and of Bm2 (all coboundaries and ggN).
[…]
$X$,Meet:=NullModN(ALS,ALC);
// <Meet> = <ALS> $\cap$ <ALC>
comp, bas, x := ComplementIntMat(ALS,Meet);
// x corresponds to the set of elements of finite order in
// <ALS> / ( <ALS> cap <ALC> ).
// Note that all these elements are of finite order.
[…]
print A_S(G) = prod_i Z/x[i]Z where ”;
[…]